OpenStudy (anonymous):

HELP PLEASE! WILL FAN AND GIVE MEDAL The function f(x) = −x2 + 48x − 432 models the profit, in dollars, a game developer makes for selling a new game, where x is the number of games sold and f(x) is the amount of profit. Part A: Determine the vertex. What does this calculation mean in the context of the problem? Part B: Determine the x-intercepts. What do these values mean in the context of the problem?

OpenStudy (jack1):

so how do u find x intercepts?

OpenStudy (jack1):

x intercepts occur where y = 0

OpenStudy (jack1):

so put f(x) = 0 into ur equation and solve for x first

OpenStudy (jack1):

its easier if u factorise the equation first tho

OpenStudy (anonymous):

I'm not 100% sure on how to factorise it completely

OpenStudy (jack1):

all good: -x^2 + 48x -432 so what 2 numbers multiply to give 432... but adds to give 48? factors of 432 = 1,432, 2,216, 3,144, 4,108 6,72 8,54 9,48 12,36 16,27 18,24 so which 2 add to give 48 answer = 12 and 36

OpenStudy (jack1):

-x^2 + 48x -432 f(x) = ( -x +36 )( x - 12)

OpenStudy (jack1):

FOIL that: ( -x +36 )( x - 12) ( -x^2 +36x + 12x -432)

OpenStudy (jack1):

so, can you solve x intercepts?: f(x) = −x2 + 48x − 432 x intercepts are when f(x) = 0 0 = −x2 + 48x − 432 0 = ( -x +36 )( x - 12) therefore x = ??? or x = ???

OpenStudy (anonymous):

I think I'm overthinking it xD

OpenStudy (anonymous):

I got some crazy numbers

OpenStudy (triciaal):

@pooja195

OpenStudy (jack1):

ok... well what about this: have u ever heard of the Quadratic Formula?

OpenStudy (anonymous):

Yes, but I haven't learned anything about it

OpenStudy (jack1):

u can use this on a calculator or write out shorthand and solve urself using the values from ur equation in ur case: ur equation is: y = ax^2 + bx + c so a = -1 b = 48 c = -432 plug in and solve

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Right?

OpenStudy (jack1):

or, easier: 0 = ( -x +36 )( x - 12) therefore either 0 = -x +36 or 0 = x - 12 ... yep, so one value is x = 12 so the other x value must be

OpenStudy (anonymous):

36

OpenStudy (jack1):

perfect! so... the hard part... what does that mean?

OpenStudy (jack1):